# Entropy conservative and entropy stable solid wall boundary conditions for the resistive magnetohydrodynamic equations

### v1

@@ Abstract
journal: Journal of Computational Physics on 13th November 2024
@@ 1 Introduction
We demonstrate the accuracy of the proposed boundary conditions through a convergence study towards an analytical solution in a circular pipe under a varying wall conductance parameter. The choice of pipe flow as a case study is strategic, showcasing the method’s efficacy in curved geometries, where meshes with curved elements are often used in conjunction with high-order accurate schemes. To verify the entropy conservative properties of the spatial discretization algorithm and our proposed hydrodynamic, thermal, and magnetic boundary conditions, we study the time rate of change of the entropy function of a flow aroud a stationary rotating spheroid. Finally, we test the new boundary conditions in two applications: fluid flow in a microchannel and a self-field magnetoplasmadynamic thruster.
@@ 2.3 Thermodynamic properties of the system
If the solution is smooth, contracting the entropy variables with the conservation equations (2.5) yields the (scalar) conservation equation for the entropy function,
@@
where the subscript $n$ denotes normal component to the wall of the corresponding tensor. Therefore, the prescribed boundary entropy flux (or entropy transfer) function $g(t)$ bounds the diffusive contribution to the time derivative of the entropy in (2.22). $\square$
@@ 3.2 Entropy conservation for internal nodes
Next, we demonstrate that the appropriate choice of SAT terms in (3.12) through (3.19) guarantees entropy dissipation at the interfaces between adjacent mesh elements. The scalar equation describing the rate of change of the entropy function in the whole domain is obtained by adding contributions from (3.11) over all the mesh elements. For any node $L$ on the interface between two neighboring mesh elements, a node $R$ exists that belongs to the adjacent element and is spatially coincident with node $L$. We will separately examine the advective and diffusive fluxes to show that for every $L$ and $R$ node pair, the total entropy production is zero for conservative fluxes and non-positive for entropy dissipative fluxes.
Terms (3.12) and (3.13) are the contributions to the time rate of change of the entropy function due to the advective SATs for the boundary nodes with indices 00 and $N$. Their cumulative contribution to $dS_{j}/dt$ is zero, as can be shown by using the generalized Tadmor’s condition (3.10):
$\frac{dS^{\text{a,cons}}_{L+R}}{dt}=\left(\mathbf{w}_{R}^{T}{\mathbf{f}}^{*}_{(R,L)}-\Psi^{*}_{R}\right)-\left(\mathbf{w}_{L}^{T}{\mathbf{f}}^{*}_{(L,R)}-\Psi^{*}_{L}\right)=\left(\mathbf{w}_{R}^{T}\mathbf{f}^{*}_{(R,L)}-\mathbf{w}_{L}^{T}\mathbf{f}^{*}_{(L,R)}\right)-\left(\Psi^{*}_{R}-\Psi^{*}_{L}\right)=0.$
@@
$\frac{dS^{\text{a,diss}}_{L+R}}{dt}=\mathbf{w}_{R}^{T}\text{DISS}^{a}_{(R,L)}+\mathbf{w}_{L}^{T}\text{DISS}^{a}_{(L,R)}=-[\![\mathbf{w}]\!]_{(L,R)}^{T}\hat{D}[\![\mathbf{w}]\!]_{(L,R)},$
@@
$\frac{dS^{\nu,\text{cons}}_{L+R}}{dt}=\left(\hat{\mathbf{w}}_{(L,R)}-\mathbf{w}_{L}\right)^{T}\mathbf{f}^{\nu}_{L}-\left(\hat{\mathbf{w}}_{(R,L)}-\mathbf{w}_{R}\right)^{T}\mathbf{f}^{\nu}_{R}+\mathbf{w}_{L}^{T}\hat{\mathbf{f}}^{\nu}_{(L,R)}-\mathbf{w}_{R}^{T}\hat{\mathbf{f}}^{\nu}_{(R,L)}=0.$ (3.21)
@@
$\frac{dS^{\nu,\text{diss}}_{L+R}}{dt}=\mathbf{w}_{L}^{T}\text{DISS}^{\nu}_{(L,R)}+\mathbf{w}_{R}^{T}\text{DISS}^{\nu}_{(R,L)}=-[\![\mathbf{w}]\!]^{T}_{(L,R)}\hat{L}[\![\mathbf{w}]\!]_{(L,R)},$
@@ 4 Entropy conservative solid wall boundary conditions
Similarly to the approach proposed in [21, 20] for the compressible Navier–Stokes equations, we must ensure that the discrete imposition of the solid wall boundary conditions for the GLM-MHD system (2.5) leads to an appropriate and physically consistent entropy production or dissipation.
@@ 4.1 Advective numerical flux divergence operator
Now, we examine the effect of the dissipation operator $\text{DISS}^{a}$ applied to the solution jump at the boundary. We consider the local Lax-Friedrichs (LLF) dissipation, modified to act in entropy space,
$\text{DISS}^{a,\text{LLF}}_{(i,j)}=\frac{1}{2}|\lambda|_{\text{max}}\underline{\mathcal{H}}[\![\mathbf{w}]\!]_{(i,j)},$ (4.3)
where $|\lambda|_{\text{max}}$ is the maximum absolute eigenvalue of the flux Jacobian considering both the $i$ and $j$ states, and the transformation matrix $\underline{\mathcal{H}}$ is a numerical analog of the continuous matrix $\mathcal{H}={\partial\mathbf{u}}/{\partial\mathbf{w}}$ presented in B. Our proof is quite general and allows for several possible definitions of $\underline{\mathcal{H}}$. As one possibility, ¯$\underline{\mathcal{H}}=\mathcal{H}(\overline{\mathbf{u}})$ can be evaluated at some average state $\overline{\mathbf{u}}=\overline{\mathbf{u}}(\mathbf{u}_{i},\mathbf{u}_{j})$. Another option, proposed in [38], defines ¯=¯$\underline{\mathcal{H}}=\underline{\mathcal{H}}(\mathbf{u}_{i},\mathbf{u}_{j})$ to be symmetric positive semi-definite and satisfying the condition $[\![\mathbf{u}]\!]=\underline{\mathcal{H}}[\![\mathbf{w}]\!]$ in all variables except energy to negotiate the difference between the LLF operators applied in conservative and entropy spaces. In both cases, the entries of $\underline{\mathcal{H}}$ result from some averaging procedure between states $i$ and $j$ denoted here with an overline. The following theorem substantiates the restriction on the average procedure, guaranteeing entropy stability of the LLF dissipation operator at the boundary.
@@
The local Lax-Friedrich dissipation operator (4.3) with the boundary state (4.1) is entropy stable at the boundary if in the definition of the matrix $\underline{\mathcal{H}}$, the average state $\overline{\mathbf{u}}$ satisfies
$\overline{v}_{n}=0,\quad\overline{B}_{n}=B^{0}_{n},\quad\overline{T}>0,\quad\overline{\rho}>0,$ (4.4)
where $\overline{v}_{n}$ and $\overline{B}_{n}$ are the normal components of the corresponding average vectors.
@@
Therefore, we represent here only the two columns of matrix $\underline{\mathcal{H}}$, interacting with the nonzero elements in (4.5). Under the assumption (4.4), the dissipative term equals
$\text{DISS}^{a,\text{LLF}}_{(-),(+)}=\frac{1}{2}|\lambda|_{\text{max}}\underline{\mathcal{H}}[\![\mathbf{w}]\!]_{(-),(+)}=\frac{1}{2}|\lambda|_{\text{max}}\begin{bmatrix}0&0\\ \overline{\rho}\overline{T}&0\\ 0&0\\ 0&0\\ 0&\overline{T}B^{0}_{1}\\ 0&\overline{T}\mu_{0}\\ 0&0\\ 0&0\\ 0&0\\ \end{bmatrix}\begin{bmatrix}-2v_{1}/T\\ 2(B^{0}_{1}-B_{1})/\mu_{0}T\end{bmatrix}.$ (4.6)
@@
$\frac{dS^{a,\text{LLF}}_{(-)}}{dt}=\mathbf{w}^{T}_{(-)}\text{DISS}^{a,\text{LLF}}_{(-),(+)}=-|\lambda|_{\text{max}}\frac{\overline{T}}{T^{2}}\left(\overline{\rho}v_{1}^{2}+\frac{(B^{0}_{1}-B_{1})^{2}}{\mu_{0}}\right).$ (4.7)
Positivity of the average density $\overline{\rho}$ and temperature $\overline{T}$ guarantees nonpositivity of the expression (4.7) and thus the entropy stability of the LLF operator. $\square$
Both the LLF dissipation operator defined in [19] or the LLF operator defined from the simple arithmetic average between two states satisfy the conditions of Theorem 4. Therefore, we attain entropy stability at the boundary and interior nodes.
@@ 4.2 Diffusive numerical flux divergence operator
We consider two different cases. The first one covers electrically insulating walls, cf. (2.25). The second one covers the electrically conducting walls, cf. (2.26). Boundary conditions for the velocity and temperature follows (2.23), (2.24). In the following theorems, we define manufactured states and gradients in the proper and general three-dimensional setting, while the proofs are carried out under the one-directional assumption for simplicity. In our notation, we omit the subscript ()(-)( - ) on the individual components of the state $\mathbf{v}_{(-)}=(\rho,\vec{v},T,\vec{B},\psi)^{T}$, and gradient $\bm{\theta}_{(-)}^{T}=(\theta_{(\rho)}^{T},\theta_{(\vec{v})}^{T},\theta_{(T)}^{T},\theta_{(\vec{B})}^{T},\theta_{(\psi)}^{T})$, where $\theta_{(\cdot)}$ refers to the gradient components corresponding to a particular set of primitive variables.
@@ 5 Numerical results
Unless otherwise stated, we define the damping parameter $\alpha^{\prime}$ as a multiple of the hyperbolic divergence cleaning speed $c_{h}^{\prime}$, in particular we use $\alpha^{\prime}=c_{h}^{\prime}/0.18$ following recommendations from [19, 26]. The value of $c_{h}^{\prime}$ is chosen on a case-by-case basis so that it does not negatively impact time step size. Unless otherwise stated, the local Lax–Friedrichs dissipation [54] and the visco-resistive dissipation from 3.22 are applied to mesh element interfaces.
@@ 5.1 Convergence study
We study three different cases for the wall conductance parameter: $c=0$, corresponding to an electrically insulating wall, $c=1$, corresponding to an electrically conducting wall, and $c=\infty$, corresponding to an electrically perfectly conducting wall. The first case tests the electrical insulating properties of the discrete wall boundary conditions (4.12), while the other two test the electric conductivity properties of the discrete wall boundary conditions (4.13). In all cases, the Hartmann number is set to $\operatorname{Ha}=5$. At this value of the Hartmann number, electromagnetic forces are strong enough to deflect the axial velocity distribution from an axisymmetric parabolic profile, while the Hartman boundary layer is still thick enough to allow for the use of relatively coarse meshes as needed in a convergence study.
@@
The results for the second, third, and fourth-order approximation of the numerical solution, corresponding to $p=1,2,3$ polynomial degrees and $N=2,3,4$ LGL nodes in each coordinate direction, are shown in Figure 3, where each column corresponds to different values of $c$. For each value of $c$, the plot shows the $L_{2}$ error of the axial velocity and axial magnetic field – these are the only two solution components deviating from a uniform value in the analytical solution. In all cases, the convergence rate reaches the “expected” value of $p+1$ (dashed lines). The various data points from Figure 3 are also available in the appendix in Table 1. The two worst numbers are $-3.87$ and $-3.93$ achieved with $p=3$ at the 4-th level of refinement for the magnetic field with conducting boundary, which can be rationalized by the super-convergence rates achieved for the same variables at the previous levels of refinement.
@@ 5.2 Entropy conservation study
Two different cases are considered to cover our proposed magnetic boundary conditions. In the first case, the spheroid is electrically insulating, while the cube walls are electrically conducting with wall conductance parameter $c=1$. In the second case, the spheroid surface is electrically conducting with $c=1$, while the cube walls are electrically insulating. In both cases, the external magnetic field on all boundaries is $\vec{B}^{0}$ as defined in (5.14). Note that the two cases considered verify each of the discrete conditions (4.12) and (4.13) in general circumstances, including: curvilinear surface with non-zero velocity, non-homogeneous wall-normal magnetic field, not-steady state solution and boundary conditions, the velocity exhibits a large discontinuity on the spheroid surface at $t=0$. We do not explicitly test for a perfectly conducting wall with $c=\infty$. This case corresponds to the homogeneous variant of (2.26) where the right-hand side is zero and is already covered by testing the magnetic boundary conditions with any positive value for wall conductance parameter, in particular, $c=1$.
@@
The numerical verification of the entropy production balance (5.16) is shown in Figure 5. The time rate of change of the total discrete entropy, $dS/dt$, is always negative, while the entropy production visco-resistive dissipation $DT$ is always positive. Furthermore, $dS/dt$ balances $DT$ up to values close to machine epsilon for IEEE double precision floating point arithmetic.

### v2

@@ Abstract
journal: Journal of Computational Physics
[…]
[label1]organization=Computer Electrical and Mathematical Science and Engineering Division (CEMSE), Extreme Computing Research Center (ECRC), King Abdullah University of Science and Technology (KAUST),city=Thuwal, postcode=23955-6900, country=Saudi Arabia
[…]
[label2]organization=Physical Science and Engineering Division (PSE), King Abdullah University of Science and Technology (KAUST),city=Thuwal, postcode=23955-6900, country=Saudi Arabia
{graphicalabstract}
@@ 1 Introduction
We demonstrate the accuracy of the proposed boundary conditions through a convergence study towards an analytical solution in a circular pipe under a varying wall conductance parameter. The choice of pipe flow as a case study is strategic, showcasing the method’s efficacy in curved geometries, where meshes with curved elements are often used in conjunction with high-order accurate schemes. To verify the entropy conservative properties of the spatial discretization algorithm and our proposed hydrodynamic, thermal, and magnetic boundary conditions, we study the time rate of change of the entropy function of a flow around a stationary rotating spheroid. Finally, we test the new boundary conditions in two applications: fluid flow in a microchannel and a self-field magnetoplasmadynamic thruster.
@@ 2.3 Thermodynamic properties of the system
If the solution is smooth, contracting the entropy variables with the balance equations (2.5) yields the (scalar) balance equation for the entropy function,
@@
where the subscript $n$ denotes the component normal to the wall of the corresponding tensor, and all the terms involving the magnetic field $\vec{B}$ cancel trivially from the vector calculus identity $\vec{C}\times(\vec{\nabla}\times\vec{D})=\vec{C}\cdot(\vec{\nabla}\vec{D}-(\vec{\nabla}\vec{D})^{T})$. Therefore, the prescribed boundary entropy flux (or entropy transfer) function $g(t)$ bounds the diffusive contribution to the time derivative of the entropy in (2.22). $\square$
@@ 3.2 Entropy conservation for internal nodes
Next, we demonstrate that the appropriate choice of SAT terms in (3.12) through (3.19) guarantees entropy dissipation at the interfaces between adjacent mesh elements. The scalar equation describing the rate of change of the entropy function in the whole domain is obtained by adding contributions from (3.11) over all the mesh elements. For any interface between two neighboring mesh elements, there exists a node $\mathcal{A}$ with index $N$ in the left mesh element and a node $\mathcal{B}$ with index 00 in the right mesh element that are spatially coincident. We will separately examine the advective and diffusive fluxes to show that for every $\mathcal{A}$ and $\mathcal{B}$ node pair, the total entropy production is zero for conservative fluxes and non-positive for entropy dissipative fluxes.
Terms (3.12) and (3.13) are the contributions to the time rate of change of the entropy function due to the advective SATs for the interface nodes $\mathcal{A}$ and $\mathcal{B}$. Their cumulative contribution to $dS/dt$ is zero, as can be shown by using the generalized Tadmor’s condition (3.10):
$\frac{dS^{\text{a,cons}}_{\mathcal{A}+\mathcal{B}}}{dt}=\left(\mathbf{w}_{\mathcal{B}}^{T}{\mathbf{f}}^{*}_{(\mathcal{B},\mathcal{A})}-\Psi^{*}_{\mathcal{B}}\right)-\left(\mathbf{w}_{\mathcal{A}}^{T}{\mathbf{f}}^{*}_{(\mathcal{A},\mathcal{B})}-\Psi^{*}_{\mathcal{A}}\right)=\left(\mathbf{w}_{\mathcal{B}}^{T}\mathbf{f}^{*}_{(\mathcal{B},\mathcal{A})}-\mathbf{w}_{\mathcal{A}}^{T}\mathbf{f}^{*}_{(\mathcal{A},\mathcal{B})}\right)-\left(\Psi^{*}_{\mathcal{B}}-\Psi^{*}_{\mathcal{A}}\right)=0.$
@@
$\frac{dS^{\text{a,diss}}_{\mathcal{A}+\mathcal{B}}}{dt}=\mathbf{w}_{\mathcal{B}}^{T}\text{DISS}^{a}_{(\mathcal{B},\mathcal{A})}+\mathbf{w}_{\mathcal{A}}^{T}\text{DISS}^{a}_{(\mathcal{A},\mathcal{B})}=-[\![\mathbf{w}]\!]_{(\mathcal{A},\mathcal{B})}^{T}\hat{D}[\![\mathbf{w}]\!]_{(\mathcal{A},\mathcal{B})},$
@@
$\frac{dS^{\nu,\text{cons}}_{\mathcal{A}+\mathcal{B}}}{dt}=\left(\hat{\mathbf{w}}_{(\mathcal{A},\mathcal{B})}-\mathbf{w}_{\mathcal{A}}\right)^{T}\mathbf{f}^{\nu}_{\mathcal{A}}-\left(\hat{\mathbf{w}}_{(\mathcal{B},\mathcal{A})}-\mathbf{w}_{\mathcal{B}}\right)^{T}\mathbf{f}^{\nu}_{\mathcal{B}}+\mathbf{w}_{\mathcal{A}}^{T}\hat{\mathbf{f}}^{\nu}_{(\mathcal{A},\mathcal{B})}-\mathbf{w}_{\mathcal{B}}^{T}\hat{\mathbf{f}}^{\nu}_{(\mathcal{B},\mathcal{A})}=0.$ (3.21)
@@
$\frac{dS^{\nu,\text{diss}}_{\mathcal{A}+\mathcal{B}}}{dt}=\mathbf{w}_{\mathcal{A}}^{T}\text{DISS}^{\nu}_{(\mathcal{A},\mathcal{B})}+\mathbf{w}_{\mathcal{B}}^{T}\text{DISS}^{\nu}_{(\mathcal{B},\mathcal{A})}=-[\![\mathbf{w}]\!]^{T}_{(\mathcal{A},\mathcal{B})}\hat{L}[\![\mathbf{w}]\!]_{(\mathcal{A},\mathcal{B})},$
@@ 4 Entropy conservative solid wall boundary conditions
Similar to the approach proposed in [21, 20] for the compressible Navier–Stokes equations, we must ensure that the discrete imposition of the solid wall boundary conditions for the GLM-MHD system (2.5) leads to an appropriate and physically consistent entropy production or dissipation.
@@ 4.1 Advective numerical flux divergence operator
Now, we examine the effect of the dissipation operator $\text{DISS}^{a}$ applied to the solution jump at the boundary. In this work, we used the local Lax-Friedrichs (LLF) dissipation, modified to act in entropy variables,
$\text{DISS}^{a,\text{LLF}}_{(i,j)}=\frac{1}{2}|\lambda|_{\text{max}}\mathcal{H}[\![\mathbf{w}]\!]_{(i,j)},$ (4.3)
where $|\lambda|_{\text{max}}$ is the maximum absolute eigenvalue of the flux Jacobian considering both the $i$ and $j$ states. The symmetric positive semi-definite matrix $\mathcal{H}=\mathcal{H}(\mathbf{u}_{i},\mathbf{u}_{j})$ is a numerical entropy Jacobian satisfying the relation
$\left([\![\mathbf{u}]\!]\right)_{i}=\left(\mathcal{H}[\![\mathbf{w}]\!]\right)_{i},\quad i=1,2,3,4,6,7,8,9,\quad\left([\![\mathbf{u}]\!]\right)_{5}\approx\left(\mathcal{H}[\![\mathbf{w}]\!]\right)_{5},$ (4.4)
following the construction recipes from [38, 19] to reconcile differences between LLF dissipation operators formulated in conservative vs. entropy variables. The expression for $\mathcal{H}$ is reproduced in Appendix B. The following theorem guarantees entropy stability of the LLF dissipation operator at the boundary.
@@
The local Lax-Friedrich dissipation operator (4.3) with the boundary state (4.1) is entropy stable at the boundary.
@@
Therefore, we represent in (4.6) only the two columns of matrix $\mathcal{H}$, interacting with the nonzero elements in (4.5). The ghost state (4.1) results into the identities $\{\!\!\{v_{1}\}\!\!\}=0$, $\{\!\!\{B_{1}\}\!\!\}=B^{0}_{1}$, and $\overline{p}=\rho RT$, simplifying expression (4.3) at the boundary node
$\text{DISS}^{a,\text{LLF}}_{(-),(+)}=\frac{1}{2}|\lambda|_{\text{max}}{\mathcal{H}}[\![\mathbf{w}]\!]_{(-),(+)}=\frac{1}{2}|\lambda|_{\text{max}}\begin{bmatrix}0&0\\ {\rho T}&0\\ 0&0\\ 0&0\\ 0&{T}B^{0}_{1}\\ 0&{T}\mu_{0}\\ 0&0\\ 0&0\\ 0&0\\ \end{bmatrix}\begin{bmatrix}-2v_{1}/T\\ 2(B^{0}_{1}-B_{1})/\mu_{0}T\end{bmatrix}{=|\lambda|_{\text{max}}\begin{bmatrix}0\\ -\rho v_{1}\\ 0\\ 0\\ B^{0}_{1}(B^{0}_{1}-B_{1})/\mu_{0}\\ (B^{0}_{1}-B_{1})\\ 0\\ 0\\ 0\\ \end{bmatrix}}.$ (4.6)
@@
$\frac{dS^{a,\text{LLF}}_{(-)}}{dt}=\mathbf{w}^{T}_{(-)}\text{DISS}^{a,\text{LLF}}_{(-),(+)}=-|\lambda|_{\text{max}}{\frac{1}{T}}\left(\rho v_{1}^{2}+\frac{(B^{0}_{1}-B_{1})^{2}}{\mu_{0}}\right).$ (4.7)
Positivity of density $\rho$ and temperature $T$ guarantees nonpositivity of the expression (4.7) and thus, the entropy stability of the LLF operator. $\square$
@@ 4.2 Diffusive numerical flux divergence operator
We consider two cases. The first one covers electrically insulating walls, cf. (2.25). The second one covers the electrically conducting walls, cf. (2.26). Boundary conditions for the velocity and temperature follows (2.23), (2.24). In the following theorems, we define manufactured states and gradients in the proper and general three-dimensional setting, while the proofs are carried out under the one-directional assumption for simplicity. In our notation, we omit the subscript ()(-)( - ) on the individual components of the state $\mathbf{v}_{(-)}=(\rho,\vec{v},T,\vec{B},\psi)^{T}$, and gradient $\bm{\theta}_{(-)}^{T}=(\theta_{(\rho)}^{T},\theta_{(\vec{v})}^{T},\theta_{(T)}^{T},\theta_{(\vec{B})}^{T},\theta_{(\psi)}^{T})$, where $\theta_{(\cdot)}$ refers to the gradient components corresponding to a particular set of primitive variables.
@@ 5 Numerical results
Unless otherwise stated, we define the damping parameter $\alpha^{\prime}$ as a multiple of the hyperbolic divergence cleaning speed $c_{h}^{\prime}$, in particular we use $\alpha^{\prime}=c_{h}^{\prime}/0.18$ following recommendations from [19, 26]. The value of $c_{h}^{\prime}$ is chosen on a case-by-case basis so that it does not negatively impact time step size. Unless otherwise stated, the modified local Lax–Friedrichs dissipation (4.3) and the visco-resistive dissipation from (3.22) are applied to mesh element interfaces.
@@ 5.1 Convergence study
We study three cases for the wall conductance parameter: $c=0$, corresponding to an electrically insulating wall, $c=1$, corresponding to an electrically conducting wall, and $c=\infty$, corresponding to an electrically perfectly conducting wall. The first case tests the electrical insulating properties of the discrete wall boundary conditions (4.12), while the other two test the electric conductivity properties of the discrete wall boundary conditions (4.13). In all cases, the Hartmann number is set to $\operatorname{Ha}=5$. At this value of the Hartmann number, electromagnetic forces are strong enough to deflect the axial velocity distribution from an axisymmetric parabolic profile, while the Hartman boundary layer is still thick enough to allow for the use of relatively coarse meshes as needed in a convergence study.
@@
The results for the second, third, and fourth-order approximation of the numerical solution, corresponding to $p=1,2,3$ polynomial degrees and $N=2,3,4$ LGL nodes in each coordinate direction, are shown in Figure 3, where each column corresponds to different values of $c$. For each value of $c$, the plot shows the $L_{2}$ error of the axial velocity and axial magnetic field – these are the only two solution components deviating from a uniform value in the analytical solution. In all cases, the convergence rate reaches the “expected” value of $p+1$ (dashed lines). The various data points from Figure 3 are also available in the appendix in Table 1.
@@ 5.2 Entropy conservation study
Two cases are considered to cover our proposed magnetic boundary conditions. In the first case, the spheroid is electrically insulating, while the cube walls are electrically conducting with wall conductance parameter $c=1$. In the second case, the spheroid surface is electrically conducting with $c=1$, while the cube walls are electrically insulating. In both cases, the external magnetic field on all boundaries is $\vec{B}^{0}$ as defined in (5.14). Note that the two cases considered verify each of the discrete conditions (4.12) and (4.13) in general circumstances, including: curvilinear surface with non-zero velocity, non-homogeneous wall-normal magnetic field, not-steady state solution and boundary conditions, the velocity exhibits a large discontinuity on the spheroid surface at $t=0$. We do not explicitly test for a perfectly conducting wall with $c=\infty$. This case corresponds to the homogeneous variant of (2.26) where the right-hand side is zero and is already covered by testing the magnetic boundary conditions with any positive value for wall conductance parameter, in particular, $c=1$.
@@
The numerical verification of the entropy production balance (5.16) is shown in Figure 5. Due to the small Reynolds numbers selected for this numerical conservation study, the time-step size is quite small ($\Delta t\approx 1.9\times 10^{-8}$). Thus, we show some results only for the first $1,300$ time-steps. The time rate of change of the total discrete entropy, $dS/dt$, is always negative, while the entropy production visco-resistive dissipation $DT$ is always positive. Furthermore, $dS/dt$ balances $DT$ up to values close to machine epsilon for IEEE double precision floating point arithmetic.